To save for retirement, you deposit $719.58 at the end of every month. If you make these payments for the next 46 years while you work, and the interest rate is 12.78%, how much will you have in your retirement account when you go to retire? That is, immeidately after the last payment is made

Answer =

Cash flow =

r =

n =
Answer #1

Balance in retirement account when you go to retire is   $ 2,35,17,463.16

Future value of cash flow = Monthly cash flow * Future value of annuity of 1
= $                 719.58 * 32682.207898
= $ 2,35,17,463.16
Working:
Future value of annuity of 1 = (((1+i)^n)-1)/i Where,
= (((1+0.010667)^552)-1)/0.010667 i = 12.8%/12 = 0.010667
= 32682.2079 n = 46*12 = 552